march correll house of couture orders and distributes high fashion clothing items to fachion outlets all over Kuala Lumpur. they buy their clothing item from a well-known designer in italy. clothing items are ordered in lots that requre quantitites of 200 units per order, and each order cost usd50 to make. demand for high fashion clothing items is 7,000 units per month and the carrying const of usd0.50 per item: i. calculate the total inventory cost if marc correll haouse of couture needs to place order per year
1
Expert's answer
2012-10-04T08:25:32-0400
total inventory cost = 12*(7000/200*50 + 0.5*7000) = 12*(1750 + 3500) = 12*5250 = $63000